OpenStudy (anonymous):

f^-1(x) = 3 root x-7 How do I figure this out?

sam (.sam.):

What you wanna figure out?

OpenStudy (anonymous):

the problem reads find f^-1(x) f(x)=3sqrroot x-7

OpenStudy (anonymous):

not sure what to do

sam (.sam.):

First let y=f(x), Think of inverse, you have to find "x" in terms of "y", \[f(x)=3\sqrt{x-7} \\ \\ y=3\sqrt{x-7} \\ \\ \frac{y}{3}=\sqrt{x-7} \\ \\ x-7=\frac{y^2}{9} \\ \\ x=\frac{y^2}{9}+7 \\ \\ f^{-1}(x)=\frac{x^2}{9}+7\]

OpenStudy (anonymous):

ok